Bonita Company owns 9,000 acres of timberland purchased in 2009 at a cost of $1,526 per acre. At the time of purchase, the land without the timber was valued at $436 per acre. In 2010, Bonita built fire lanes and roads, with a life of 30 years, at a cost of $91,560. Every year, Bonitasprays to prevent disease at a cost of $3,270 per year and spends $7,630 to maintain the fire lanes and roads. During 2011, Bonita selectively logged and sold 763,000 board feet of timber, of the estimated 3,815,000 board feet. In 2012, Bonita planted new seedlings to replace the trees cut at a cost of $109,000.

Determine the depreciation expense and the cost of timber sold related to depletion for 2011. (Round intermediate calculations to 5 decimal places, e.g. 1.54687 and final answers to 0 decimal places, e.g. 5,125.)

Depreciation expense

$

Cost of timber sold

$

  

  

Bonita has not logged since 2011. If Bonita logged and sold 981,000 board feet of timber in 2019, when the timber cruise (appraiser) estimated 5,450,000 board feet, determine the cost of timber sold related to depletion for 2022. (Round intermediate calculations to 5 decimal places, e.g. 1.54687 and final answers to 0 decimal places, e.g. 5,125.)

Cost of timber sold

$
Answer #1

A. Determine the depreciation expense and the cost of timber sold related to depletion for 2011.

Cost of Timberland = 1526

Less:Cost of Land = 436

Cost of Timber = 1090

Value of Timber = 9,810,000 (9000*1090)

Estimated board fee = 3815000

Total cost per board ft. = 2.57142 (9810000/3815000)

Board feet timber sold = 763000

Cost of timber sold to depletion = 2.57142*763000 = 1961993

Fire lanes and road Life 30yrs

Fire lanes and road Cost 91560

Depreciation expense = 3052 = (91,560/30)

B.Cost of timber sold = 9810000-1961993 - (7630+3270)*30 + 109000 = 7630007

Cost of timber sold = 7630007/5450000*981000 = 1373401